The conjugate of ammonia NH4+, is a weak acid. If a 0.200 M NH4Cl solution has a pH of 5.0, what is the Ka of NH4+?
1
Expert's answer
2020-01-28T06:16:00-0500

NH4Cl dissociates as:

NH4Cl <--> NH4+ + Cl-

From here:

Ka = ([NH4+][Cl-])/[NH4Cl] = [NH4+]2/[NH4Cl]

As pH = -log[H+]:

[H+] = 10-pH

From here, the concentration of [NH4+] in the solution equals:

[H+] = 10-5 M

As [H+] = [NH4+]= 10-5 M, [NH4Cl] = 0.200 M, Ka of NH4+ equals:

Ka = [NH4+]2/[NH4Cl] = (10-5)2 / 0.2 = 5 × 10-10


Answer: 5 × 10-10